The Film Company uses a job-order costing system. The following information pertains to the company's activities for the month of December:

On December 1, the company had two jobs in work in process as follows: Job 6 $20,000 and Job 8 $14,000

On December 1, the company had job in Finished Goods as follows: Job 10 $50,000

Jobs 12 and 13 were started during the month.

Materials were purchased on account in December for $54,000

Materials were requisitioned for use in production. The direct materials were distributed as follows:

Job #6 $6,000 Job #8 $8,000 Job #12 $23,000 Job #13 $7,000

Direct labor was incurred in December. The direct labor was distributed as follows:

Job #6 $30,000 Job #8 $20,000 Job #12 $35,000 Job #13 $10,000

Overhead is applied using camera hours. The budgeted overhead for the coming year is $1,000,000

Estimated camera hours is the cost driver and is 250,000 camera hours

The following is the distribution of Camera hours (CH) for December:

Job #6 3,000 CH Job #8 4,000 CH Job #12 11,000 CH Job #13 2,000 CH

Actual overhead for December was incurred for various indirect costs on account.  

Use a generic miscellaneous payable account in your journal entry. $75,000

Job Nos. 8 and 12 were completed during the month.

Job Nos.10 and 12 were sold to customers on account during December at 125% of cost

Any over-or underapplied overhead is closed to Cost of Goods Sold.

Questions

1. What was the total beginning balance (December 1st) of Work in process?

2. What was the beginning balance (December 1st) of Finished Goods?

3. Calculate the overhead rate based on camera hours
Answer #1

Answer 1:

Total beginning balance (December 1st) of Work in process = Work in process of Job 6 + Work in process of Job 8

= 20000 + 14000

= $34,000

Total beginning balance (December 1st) of Work in process = $34,000

Answer 2:

Beginning balance (December 1st) of Finished Goods = Job 10 Finished Goods = $50,000

Beginning balance (December 1st) of Finished Goods = $50,000

Answer 3:

Overhead rate based on camera hours = budgeted overhead / Estimated camera hours = 1000000 / 250000 = $4

Overhead rate based on camera hours = $4 per camera hour